OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1

OpenStudy (anonymous):

Help me plz!!

sam (.sam.):

y=9

sam (.sam.):

5x + 4y = 1 5x = 1-4y x=(1-4y)/5 4x + 3y = -1 4((1-4y)/5) + 3y = -1 y=9

OpenStudy (anonymous):

5x + 4y = 1 4x + 3y = -1 -> ( -7, 9)